OpenStudy (anonymous):

t = 2π √l/g Re-arrange to make l the subject.

OpenStudy (xapproachesinfinity):

in other words solve for l (this by the way something in physics, but can't remember what it is) \(t=2\pi\sqrt{\frac{l}{g}}\\\frac{t}{2\pi}=\sqrt{\frac{l}{g}}\) square both sides \((\frac{t}{2\pi})^2=(\sqrt{\frac{l}{g}}~)^2\)

OpenStudy (xapproachesinfinity):

continue with the rest!

OpenStudy (anonymous):

thanks

OpenStudy (xapproachesinfinity):

welcome